Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote mode.

Mar 16th, Tuesday

16h - 21h; Angel, Bilal: Smooth start of the pool, no problem to report, mediocre weather starting but improving a lot during the evening.

21h - 03h; Golshan and Suzanne: We started out a bit blocked when the desktop screens froze and had to be restarted (our restarting VNC did not solve the problem). We could not change desktops as usual and Bilal found a solution. The tau values were always < 0.18 and most of the night~ 0.15. Early on the sky fluctuations were not great (5 to 10 Jy/beam on array2), but improved during the night. We did a beam map on 3C279. We were not able to do a calib scan on CRL618 as it was too low. At the end of the shift we did a skydip We observed 160-16 NGC3627 (IMEGIN) for 4 repetitions (2h) this galaxy is not yet completed - needs 4 more hours to complete. We observed 180-20 project for 2 hours.

Mar 17th, Wednesday

03h - 09h; Isabelle and Karine: The observations conditions were excellent : with an opacity of _0.15, very stable, and perfect beams. We first observed for project 181-20 , source RXJ1347 , with 2 Nreps (2h), we then observed for project 107-20 on L1689A with 2 Nreps (~ 2h), then we observed GP_L23p3 from GASTON Galactic plane project, with 2 Nreps (1h).

09h - 15h; Chenlin: The opacity was good (around 0.15). We started the observations for project 122-16 with one repetition of GP_L23p9 and two repetitions of GP_L24p5. Then we observed for project 175-20 with one repetition of NEP_L. From 2pm, the beam size became larger and asymmetric. We did calib_1scan and a skydip. And we kept doing pointing and focus to monitor the beam till the end of the shift.

15h - 21h; Bilal: The sky conditions and beam parameters were bad, but improved a lot by 5pm.

21h - 03h; Golshan and Suzanne: It was a very stable night: The tau values were always < 0.2, hovering around ~0.18 most of the night. The sky fluctuations were quite good and steady all night: ~ +/-5 to 3 Jy/beam on array2. We did a calib1 scan on CRL618. A beam map on 3C279 was started at the end of our shift. We observed 160-16 NGC3627 (IMEGIN) for 4 repetitions (2h). Then we observed project 180-20 for 2 hours until it dropped to 30 deg elevation (the limit as stated in the readme file is 35 to 30 deg min). We returned to NGC3627 for 160-16 for the last hour - it was very close by to the 180-20 source.

Mar 18th, Thursday

03h - 09h; We first made a pointing + focus + beammap on 3C279 cat NIKA2R6; Then we have observed for project 181-20 : source RXJ1347 with 1 Nrep, then for project 107-20 : source L1689A with 1 Nrep. Then, at ~5h30 we switched to GASTON_dust project, with 1 Nrep on L1689B, and GASTON_GP with 2 Nrep on L23p9 and started 1 Nrep on L24p5.

09h - 15h; Chenlin: Excellent opacity, continuing on GASTON_GP, then 175-20 project and then 095-20. Acquisition crashed around 11AM LT. It was rebooted, one or two paintings affected.

15h - 21h; Bilal: Continuing on 095-20, but the atmosphere is a bit unstable. Degrading fast in the end of the afternoon with very thick cloud coverage.

21h - 03h; Suzanne, Bilal: Unstable beam and weather, and conditions degrading slowly until snow arrived.

Mar 19th, Friday

03h - 09h; Aurore: Stopped part of the night, starting again around 5AM in surprisingly excellent conditions.

09h - 15h; Anika: Excellent conditions.

15h - 21h; Bilal: Observations stopped part of the afternoon, and starting again in improving conditions.

21h - 03h; Longji: Smooth operation.

Mar 20th, Saturday

03h - 09h; Andrew: Dish has been parked due to snow, but is ready to go as I take over. Beams are really excellent and tau < 0.2. Observed one repeat of GOODSNORTH for N2CLS, followed by 2 repeats of L1689A for 107-20 with tau~0.12. Snow stops observations as I am about to begin observing for GASTON. Naturally...

09h - 15h; Anika: Starting and stopping through the morning.

15h - 21h; Ankur: At the beginning of the shift it started to snow and telescope was stopped. Restarted around 19.30h CET and started 1 repetition on project 180-20. Tau ~ 0.18, bit unstable +- 10Jy/Beam on array 2 - beams slightly broad.

21h - 03h; Nico N2CL2. Let the previous obs on 181-20 run, then did a focus pointing and then integrated on COSMOS. Cross checked the focus and pointing after about two hours of obs. Everything went smoothly.

Mar 21st, Sunday

03h - 09h; Andrew: A few scans finished on COSMOS. One repeat on RXJ1347 for 181-20, followed by one repeat on L1689A for 107-20. The rest of the shift was spent on the 122-16/GASTON GP fields. Conditions were excellent throughout.

09h - 15h; Observing in good conditions.

15h - 21h; We started with project 095-20 with a tau ~ 0.25, and completed 4 repetitions on source 1390. Then completed 3 repetitions on source 869. Opacity increased above 0.3, but was ~ 0.25 in evening . 2mm beams were good. Smooth run.

21h - 03h; Stable night observing NGC3627 for 160-16, and COSMOS for the rest of the shift.

Mar 22nd, Monday

03h - 09h; Andrew: Finished a beam map as I took over. Acquired two more repeats on RXJ1347 for 181-20, then spent the rest of the time on the 122-16/GASTON GP fields. Conditions were fantastic, with extremely stable beams. I expect we might finish this part of GASTON today or tomorrow.

09h - 15h; No problem in the morning, observing in good opacity.

15h - 21h; Continuing observations.

21h - 03h; Smooth night on COSMOS, good focus session. La quille !

Mar 23rd, Tuesday

03h - 05h; End of the pool on 181-20 under excellent conditions.

And that's all folks!

DailyReportsNika2Pool160321 (last edited 2021-03-23 12:49:20 by NikaBolometer)